OpenStudy (anonymous):

hat is the value of the expression when a = 4 and b = 2? A. 48 B. 80 C. 120 D. 136

OpenStudy (anonymous):

@uri

OpenStudy (uri):

where is the expression? @RasheedahOwl

OpenStudy (anonymous):

2a^3 + 2b^2

OpenStudy (uri):

plug in 4 instead of a and 2 instead of b,can you show me? @RasheedahOwl

OpenStudy (anonymous):

k

OpenStudy (anonymous):

2(4)^3+2(2)^2

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

is the answer 136?

OpenStudy (uri):

Yeah,Now solve the exponents first 4^3=4*4*4= 2^2=2*2=

OpenStudy (uri):

I didn't check. Let's solve it and we'll get to the answer.

OpenStudy (anonymous):

k.

OpenStudy (anonymous):

4^3=64

OpenStudy (anonymous):

then. 64*2

OpenStudy (anonymous):

=128

OpenStudy (anonymous):

then. 2^2=4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4*64=128 + 2*4=8

OpenStudy (anonymous):

All = 136

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

soryy told u wrong

OpenStudy (anonymous):

i ment 64*2=128

OpenStudy (uri):

yes the answer is 136.

OpenStudy (uri):

We got confused with 4 instead of 2,silly mistake ha anyways. :)